# Vision-EKIPL: External Knowledge-Infused Policy Learning for Visual Reasoning

### v1

@@ 1 Introduction
To effectively enhance the visual reasoning capability of machines, the research community has explored diverse technical approaches. Current mainstream research paradigms can be broadly categorized into three types: (1) neural-symbolic methods [13, 4, 8, 67, 16], which aim to integrate the exceptional pattern recognition strengths of deep neural networks with the inherent logical rigor and interpretability of symbolic systems. (2) Supervised Fine-Tuning (SFT) of Multimodal Large Language Models(MLLMs) [58, 54], which relies on large-scale annotated datasets for end-to-end training to directly optimize model performance on specific visual reasoning tasks. (3) Reinforcement Learning (RL) based methods [52, 24], exemplified by techniques (e.g., Group Relative Policy Optimization (GRPO) [47]). Such methods leverage RL’s reward mechanisms to guide and activate the latent reasoning potential within pretrained foundation models, demonstrating favorable generalization capability, particularly when tackling complex visual-cognitive tasks involving mathematical logic derivation or code understanding, thus garnering increasing attention.
@@
Figure 1: The output examples of Vision-EKIPL on three visual reasoning tasks.
@@ 2 Related Work
Visual reasoning is a key research direction in artificial intelligence, aiming to enable machines to understand and interpret visual information and, on that basis, perform complex cognitive tasks such as logical inference, causal analysis, and problem solving. This technology has broad application prospects, including visual counting [38, 37], geometric problem solving [12, 31, 41, 68, 48], visual transformation reasoning [22], scientific research [40, 32], and robotic task planning [23, 29, 17]. Early work in visual reasoning relied on programmatic generation [30, 16, 51] or neuro-symbolic methods [13, 4, 8, 67]. In recent years, driven by the rapid development of Multimodal Large Language Models(MLLMs), the field has seen breakthrough progress. For example, LLaVA-CoT [58] employs a multi-stage Chain-of-Thought (CoT) [57] supervised fine-tuning (SFT) strategy, while Insight-V [11] combines SFT with reinforcement learning (RL). DeepSeek-R1-Zero [14] introduced a rule-based RL approach, significantly enhancing reasoning capability. Building upon DeepSeek-R1 [14], we propose a novel RL method that substantially improves the model’s reasoning performance.
@@
Reinforcement learning (RL) has demonstrated significant efficacy in enhancing the reasoning capabilities of Large Language Models (LLMs) through iterative, feedback-driven refinement [9, 49, 47, 60, 62, 26, 69]. Notable methodologies include Reinforcement Learning from Human Feedback (RLHF) [44] and Reinforcement Learning from AI Feedback (RLAIF) [6], both of which leverage either human or AI-generated feedback to refine model behavior. Within the domain of vision-language tasks, RL has been successfully employed to align model predictions with human preferences and mitigate the occurrence of hallucinations [50, 63, 64, 70]. More recently, advancements such as DeepSeek-R1-Zero [15] have introduced Group Relative Policy Optimization (GRPO) [47], a technique that utilizes rule-based rewards to strengthen reasoning abilities without requiring supervised fine-tuning. GRPO has been further adapted for specialized applications, with Visual-RFT [39] employing it for visual grounding and Med-R1 [45] applying it to medical reasoning tasks.Vision-R1 [24] and Reason-RFT [52] adopt a two-stage training paradigm—CoT supervised fine-tuning followed by GRPO-based reinforcement fine-tuning—to enhance the reasoning performance of Multimodal Large Language Models (MLLMs). Distinctly, our Vision-EKIPL is the first to leverage high-quality actions generated by external models to guide policy-model optimization, thereby infusing novel reasoning knowledge and pushing forward the boundary of model’s reasoning capacity.
@@ 3.1 Preliminaries
The policy model parameters $\theta$ are then updated to increase the likelihood of selecting actions that demonstrated positive advantages and decrease the probability of choosing actions with negative advantages. This update is typically performed using gradient-based optimization method. To mitigate excessive policy updates and enhance training stability, GRPO constrains $\text{ratio}_{i}$ to the interval $[1-\delta,1+\delta]$. Moreover, to encourage the learned policy to remain in proximity to the reference distribution $\pi_{\text{ref}}$, a Kullback-Leibler (KL) divergence penalty, weighted by a coefficient $\beta$, is integrated into the optimization objective. Finally, the optimization objective of GRPO can be formulated as follows:
@@ 3.2 External Knowledge-Infused Policy Learning
Sampling Action Groups beyond policy model. We introduce a total of $M$ auxiliary models to support the learning process. Given an input state $s=(x,q)$, where $x$ denotes the visual encoding of the input image and $q$ represents the textual encoding of the question, GRPO first samples a group of actions $\{o^{\theta}_{1},\dots,o^{\theta}_{N}\}$ from the current policy $\pi_{\theta}$. Additionally, for each auxiliary model $\pi^{\phi_{j}}$, it samples a corresponding group of actions $\{o^{\phi_{j}}_{1},\dots,o^{\phi_{j}}_{N}\}$. The sampling process is as follows:
@@
$o^{\phi_{j}}_{i}\sim\pi^{\phi_{j}}(o\mid x,q),\quad\text{for }i=1,2,\dots,N,\quad\text{for }j=1,2,\dots,M$ (5)
@@
  • If the response provides a correct final answer,the model receives an accuracy reward of +1. Otherwise, the model receives 0 reward.
@@
Algorithm 1 Vision-EKIPL: External Knowledge-Infused Policy Learning for Visual Reasoning
Input initial policy model $\pi_{\theta_{\text{init}}}$; reward function $R$; input states $\mathcal{S}$; external models$\{\pi^{\phi_{j}}\}_{j=1}^{M}$
1:policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
2:for iteration = 1, …, I do
3:    reference model $\pi_{ref}\leftarrow\pi_{\theta}$
4:    for step = 1, …, B do
5:         Sample a batch $\mathcal{S}_{b}$ from $\mathcal{S}$
6:         Update the old policy model $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
7:         Sample $N$ actions $\{o_{i}^{\theta}\}_{i=1}^{N}\sim\pi_{\theta_{old}}(o\mid s)$ for each input state $s\in\mathcal{S}_{b}$
8:         for j = 1,…,M do
9:             Sample $N$ actions $\{o_{i}^{\phi_{j}}\}_{i=1}^{N}\sim\pi_{\theta}^{\phi_{j}}(o\mid s)$ for each input state $s\in\mathcal{S}_{b}$
10:         Action set $O\leftarrow\{o^{\theta}_{i}\mid i=1,\dots,N\}\;\cup\;\bigcup_{j=1}^{M}\{o^{\phi_{j}}_{i}\mid i=1,\dots,N\}$
11:         Compute rewards for each sampled action within $O$ by running $R$
12:         Select the top-G reward actions from $O$ to form high-quality action set $T$
13:         Compute $A_{i}$ for the action $o_{i}$ within $T$ through group relative advantage estimation (Eqn. 1)
14:         for GRPO iteration = 1, …, $\mu$ do
15:             Update the policy model $\pi_{\theta}$ by maximizing the GRPO objective (Eqn. 3)
Output $\pi_{\theta}$
@@ 4.1 Experimental Details
In this paper, we employ the Reason-RFT-CoT Dataset [52] to evaluate our method. The experiments are organized into the following three task categories: (1) Visual Counting This task assesses multimodal reasoning by integrating linguistic, visual, and mathematical skills to solve arithmetic problems within 3D block-based scenes. (2) Structure Perception This visual reasoning task requires models to interpret structural information across various mathematical geometries, medical imaging, chart layouts, and architectural designs. (3) Spatial Transformation This spatial-visual reasoning task evaluates a model’s ability to infer single-step or multi-step transformation actions by analyzing initial and final visual states of 3D scenes presented from multiple perspectives (e.g., center, left, right). Each task contains in-domain test set and out-of-domain test set. Specific information can be found in [52]
Evaluation Metrics The primary evaluation metric is the accuracy rate (Acc) [66]. For numerical answers, correctness is determined by verifying mathematical equivalence to the ground truth. In the case of multiple-choice questions, a string match is performed against the correct option. For function-type sequences, a stepwise multi-level evaluation approach is utilized to assess alignment with the correct solution sequence.
Implementation Details In our experiments, we utilize Qwen2-VL-2B and Qwen2-VL-7B [56] as policy models. For external models, we selected GPT-4o [27] and Gemini-1.5-Pro [53]. Our implementation is based on the open-source frameworks Open-R1 [25] and vLLM [33] to ensure reproducibility of results and system scalability. All experiments were conducted on a server equipped with 8 A100 GPUs.
Baselines for comparison To evaluate the performance and generalization capabilities of different training strategies, adhering to the settings in [52], the methods compared in this paper are as follows: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with chain-of-thought (CoT) reasoning. (2) RL-based methods——Reason-RFT-Zero, which applies RL traing directly to the base model,Reason-RFT, which first performs supervised learning with partial chain-of-thought (CoT) data before RL training and Vision-EKIPL, which integrates the external auxiliary models into the RL training.
@@
$+$ ANS-SFT[52] 96.20 39.20 67.70 51.34 22.50 36.92 77.39 49.24 50.33 58.99
$+$ CoT-SFT[52] 85.50 46.50 66.00 43.05 25.25 34.15 64.37 43.19 42.86 50.14
$+$ Reason-RFT-Zero[52] 98.40 44.80 71.60 47.68 32.50 40.09 42.13 34.07 33.41 33.74
$+$ Reason-RFT[52] 96.80 51.20 74.00 49.03 33.13 41.08 74.61 64.05 64.08 67.58
$+$ Vision-EKIPL(Ours) 99.10 52.30 75.70 49.70 34.50 42.10 78.23 65.12 65.45 69.60
@@
$+$ ANS-SFT[52] 95.00 33.90 64.45 51.34 25.38 38.36 82.19 54.29 54.83 63.77
$+$ CoT-SFT[52] 87.30 42.40 64.85 50.49 33.00 41.75 81.31 47.90 47.80 59.00
$+$ Reason-RFT-Zero[52] 99.40 53.00 76.20 55.00 54.75 54.88 67.67 57.20 56.15 56.68
$+$ Reason-RFT[52] 95.60 51.00 73.30 59.27 49.25 54.26 79.97 59.36 58.61 65.98
$+$ Vision-EKIPL(Ours) 99.70 53.30 76.50 60.10 56.75 58.42 83.32 62.35 60.47 68.71
@@ 4.4 Analysis on the Sources of Action
Figure 5: Pass@k curves of base model,Reason-RFT and Vision-EKIPL on the Spatial Transformation task. (a) Evaluation results for 2B model on ID task, (b) Evaluation results for 7B model on ID task.
@@ 4.5 Push forward the boundary of the model’s reasoning ability
As shown in Fig. 5, we conducted a comparative analysis of the Pass@$K$ scores for Vision-EKIPL and Reason-RFT under different $k$ values. The Pass@$K$ metric measures the probability that the model produces at least one correct answer within $K$ independent samples, which effectively reflects the upper boundary of the model’s reasoning capability achievable through broader exploration. By providing the model with numerous opportunities to attempt to solve problems (i.e., setting a large $K$ value), we can more comprehensively evaluate the reasoning boundaries reachable by the base model as well as models trained with different RL methods. This evaluation method provides a critical and rigorous perspective for determining whether RL training genuinely enhances the model’s fundamental reasoning ability.
At smaller $k$ values, both Vision-EKIPL and Reason-RFT outperform the base model, which was not trained with any RL. However, a notable phenomenon is that as the value of $k$ continues to increase, the Pass@$K$ score of Reason-RFT begins to drop below that of the base model. This result suggests that the base model itself, through sufficiently diverse sampling exploration, possesses the ability to generate correct answers for problems traditionally thought to require RL training. This further indicates that traditional RL training may not effectively enhance, and could even to some extent limit, the model’s inherent potential reasoning scope that can be uncovered through extensive exploration.
In stark contrast, at larger $k$ values, Vision-EKIPL’s Pass@$K$ score significantly surpasses both the base model and Reason-RFT. This observation strongly highlights Vision-EKIPL’s significant potential: through the large-scale sampling facilitated by its design, Vision-EKIPL can uncover a much wider range of correct reasoning paths that are difficult for the base model and Reason-RFT to cover, thereby substantially expanding the model’s reasoning boundary. Vision-EKIPL’s ability to break through the base model’s reasoning boundary lies primarily in its core mechanism: skillfully incorporating high-quality actions generated by external models as distillation targets during the policy model update process. This innovative mechanism not only significantly enhances the model’s sampling efficiency but, more crucially, effectively infuses novel reasoning knowledge and provides a richer decision-making perspective. When exploring the space of complex problems, this enables Vision-EKIPL to demonstrate exceptional efficacy, ultimately allowing it to successfully identify and utilize correct reasoning paths that are difficult for the base model to access, thereby raising the model’s reasoning boundary.

### v2

@@ 1 Introduction
Figure 1: The output examples of Vision-EKIPL on three visual reasoning tasks, which shows superior reasoning results.
To effectively enhance the visual reasoning capability of machines, the research community has explored diverse technical approaches. Current mainstream research paradigms can be broadly categorized into three types: (1) neural-symbolic methods [13, 4, 8, 67, 16], which aim to integrate the exceptional pattern recognition strengths of deep neural networks with the inherent logical rigor and interpretability of symbolic systems. (2) Supervised Fine-Tuning (SFT) of MLLMs [58, 54], which relies on large-scale annotated datasets for end-to-end training to directly optimize model performance on specific visual reasoning tasks. (3) Reinforcement Learning (RL) based methods [52, 24], exemplified by techniques (e.g., Group Relative Policy Optimization (GRPO) [47]). Such methods leverage RL’s reward mechanisms to activate the latent reasoning potential within pretrained models, demonstrating favorable generalization capability, particularly when tackling complex visual-cognitive tasks involving mathematical logic derivation or code understanding, thus garnering increasing attention.
@@ 2 Related Work
This technology has broad application prospects, including visual counting [38, 37], geometric problem solving [12, 31, 41, 68, 48], visual transformation reasoning [22], scientific research [40, 32], and robotic task planning [23, 29, 17]. Early work in visual reasoning relied on programmatic generation [30, 16, 51] or neuro-symbolic methods [13, 4, 8, 67]. In recent years, driven by the rapid development of MLLMs, the field has seen breakthrough progress. For example, LLaVA-CoT [58] employs a multi-stage Chain-of-Thought (CoT) [57] supervised fine-tuning (SFT) strategy, while Insight-V [11] combines SFT with reinforcement learning (RL). DeepSeek-R1-Zero [14] introduced a rule-based RL approach, significantly enhancing reasoning capability. Building upon DeepSeek-R1 [14], we propose a novel RL method that substantially improves the model’s reasoning performance.
@@
Reinforcement learning (RL) has demonstrated significant efficacy in enhancing the reasoning capabilities of Large Language Models (LLMs) through iterative, feedback-driven refinement [9, 49, 47, 60, 62, 26, 69]. Notable methodologies include Reinforcement Learning from Human Feedback (RLHF) [44] and Reinforcement Learning from AI Feedback (RLAIF) [6], both of which leverage either human or AI-generated feedback to refine model behavior. Within the domain of vision-language tasks, RL has been successfully employed to align model predictions with human preferences and mitigate the occurrence of hallucinations [50, 63, 64, 70]. More recently, advancements such as DeepSeek-R1-Zero [15] have introduced GRPO [47], a technique that utilizes rule-based rewards to strengthen reasoning abilities without requiring supervised fine-tuning. GRPO has been further adapted for specialized applications, with Visual-RFT [39] employing it for visual grounding and Med-R1 [45] applying it to medical reasoning tasks.Vision-R1 [24] and Reason-RFT [52] adopt a two-stage training paradigm—CoT supervised fine-tuning followed by GRPO-based reinforcement fine-tuning—to enhance the reasoning performance of MLLMs. Distinctly, our Vision-EKIPL is the first to leverage high-quality actions generated by external models to guide policy-model optimization, thereby infusing novel reasoning knowledge and significantly advancing reasoning ability.
@@ 3.1 Preliminaries
The policy model parameters $\theta$ are then updated to increase the likelihood of selecting actions that demonstrated positive advantages and decrease the probability of choosing actions with negative advantages. This update is typically performed using gradient-based optimization method. To mitigate excessive policy updates and enhance training stability, GRPO constrains $\text{ratio}_{i}$ to the interval $1-\delta,1+\delta]$. Moreover, to encourage the learned policy to remain in proximity to the reference distribution $\pi_{\text{ref}}$, a Kullback-Leibler (KL) divergence penalty, weighted by a coefficient $\beta$, is integrated into the optimization objective. Finally, the optimization objective of GRPO can be formulated as follows:
@@ 3.2 External Knowledge-Infused Policy Learning
Sampling Action Groups beyond policy model. We introduce a total of $M$ auxiliary models to support the learning process. Although the auxiliary actions are not solely derived from $\pi_{\theta_{\text{old}}}$, as long as the set of actions drawn, for example, from previous policy iterations or a dedicated exploration space, collectively encompasses the entire trajectory space that the current target policy $\pi_{\theta_{\text{new}}}$ can generate, then employing these mixed actions in importance sampling remains theoretically valid, with unbiasedness guaranteed. We have proven it theoretically in the appendix A. Given an input state $s=(x,q)$, where $x$ denotes the visual encoding of the input image and $q$ represents the textual encoding of the question, GRPO first samples a group of actions $\{o^{\theta}_{1},\dots,o^{\theta}_{N}\}$ from the current policy $\pi_{\theta}$. Additionally, for each auxiliary model $\pi^{\phi_{j}}$, it samples a corresponding group of actions $\{o^{\phi_{j}}_{1},\dots,o^{\phi_{j}}_{N}\}$. The sampling process is as follows:
@@
$\begin{split}o^{\phi_{j}}_{i}\sim\pi^{\phi_{j}}(o\mid x,q),\quad\text{for }i=1,2,\dots,N,\\ \quad\text{for }j=1,2,\dots,M\end{split}$ (5)
@@
  • If the response provides a correct final answer,the model receives an positive accuracy reward. Otherwise, the model receives 0 reward. For the specific definition of accuracy reward, please refer to [52].
@@ 4.1 Experimental Details
In this paper, we employ the Reason-RFT-CoT Dataset [52] to evaluate our method. The experiments are organized into the following three task categories: (1) Visual Counting This task assesses multimodal reasoning by integrating linguistic, visual, and mathematical skills to solve arithmetic problems within 3D block-based scenes. (2) Structure Perception This visual reasoning task requires models to interpret structural information across various mathematical geometries, medical imaging, chart layouts, and architectural designs. (3) Spatial Transformation This spatial-visual reasoning task evaluates a model’s ability to infer single-step or multi-step transformation actions by analyzing initial and final visual states of 3D scenes presented from multiple perspectives (e.g., center, left, right). Each task contains in-domain test set and out-of-domain test set. Specific information can be found in [52].
Implementation Details In our experiments, we utilize Qwen2-VL-2B and Qwen2-VL-7B [56] as policy models. For external models, we selected GPT-4o [27] and Gemini-1.5-Pro [53]. Our implementation is based on the open-source frameworks Open-R1 [25] and vLLM [33] to ensure reproducibility of results and system scalability. For hyperparameters, we employed a cosine learning rate schedule with a peak value of $5\times 10^{-7}$ and adopted the AdamW optimizer to optimize the policy model, setting N and G to 8, with a default KL penalty of $\beta=0.005$. Based on the dataset size, we set the number of epochs for Visual Counting task, Spatial Transformation task, and Structure Perception task to 1, 1, and 5, respectively. To ensure stability and statistical significance of the results, we repeat each major experiment three times under the same setting and report the average accuracy across runs as the final result. All experiments are conducted on a server equipped with 8 A100 GPUs.
API Cost Considerations Incorporating high-quality actions generated by external models such as GPT-4o and Gemini-1.5-Pro introduces non-negligible API costs. These costs include both inference latency (typically 1–3 seconds per query depending on sequence length and model load) and monetary expenses (approximately 0.010.01–0.03 per query, varying with provider pricing tiers). Although these costs are relatively minor in small-scale experiments, they may accumulate significantly in large-scale training or deployment scenarios. In future work, exploring cost-effective open-source alternatives or distillation-based approaches to mitigate reliance on expensive APIs could be a promising direction.
Baselines for comparison To evaluate the performance and generalization capabilities of different training strategies, adhering to the settings in [52], the methods compared in this paper are as follows: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with chain-of-thought (CoT) reasoning. (2) RL-based methods——Reason-RFT-Zero, which applies RL training directly to the base model, Reason-RFT, which first performs supervised learning with partial chain-of-thought (CoT) data before RL training and Vision-EKIPL, which integrates the external auxiliary models into the RL training.
@@
++ ANS-SFT[52] 96.20 39.20 67.70 51.34 22.50 36.92 77.39 49.24 50.33 58.99
++ CoT-SFT[52] 85.50 46.50 66.00 43.05 25.25 34.15 64.37 43.19 42.86 50.14
++ Reason-RFT-Zero[52] 98.40 44.80 71.60 47.68 32.50 40.09 42.13 34.07 33.41 33.74
++ Reason-RFT[52] 96.80 51.20 74.00 49.03 33.13 41.08 74.61 64.05 64.08 67.58
++ Vision-EKIPL(Ours) 99.10 52.30 75.70 49.70 34.50 42.10 78.23 65.12 65.45 69.60
@@
++ ANS-SFT[52] 95.00 33.90 64.45 51.34 25.38 38.36 82.19 54.29 54.83 63.77
++ CoT-SFT[52] 87.30 42.40 64.85 50.49 33.00 41.75 81.31 47.90 47.80 59.00
++ Reason-RFT-Zero[52] 99.40 53.00 76.20 55.00 54.75 54.88 67.67 57.20 56.15 56.68
++ Reason-RFT[52] 95.60 51.00 73.30 59.27 49.25 54.26 79.97 59.36 58.61 65.98
++ Vision-EKIPL(Ours) 99.70 53.30 76.50 60.10 56.75 58.42 83.32 62.35 60.47 68.71
@@ 4.4 Analysis on the Sources of Action
Figure 5: Pass@k curves of base model,Reason-RFT and Vision-EKIPL on the Spatial Transformation task. Evaluation results for 2B model (a) and 7B model (b) on ID task.
@@ 4.5 Push Forward the Boundary of the Model’s Reasoning Ability
As shown in Fig. 5, we conduct a comparative analysis of the Pass@$K$ scores for Vision-EKIPL, Reason-RFT, and the base model under varying $k$ values. The Pass@$K$ metric reflects the likelihood that a model produces at least one correct answer across $K$ independent samples, and thus serves as a proxy for evaluating the upper bound of the model’s reasoning capability when given sufficient exploration opportunities.
At lower values of $k$, both Vision-EKIPL and Reason-RFT outperform the base model, indicating that reinforcement learning helps guide the model toward more accurate reasoning paths in the early stages. However, a noteworthy phenomenon emerges as $k$ increases: the Pass@$K$ score of Reason-RFT begins to fall behind that of the base model. This suggests that although Reason-RFT improves sample efficiency, it may limit the diversity of the model’s reasoning space by overemphasizing high-reward but familiar reasoning patterns, thus constraining its exploratory capacity.
In contrast, Vision-EKIPL demonstrates a consistently superior performance across all $k$ values. Especially at larger $k$, its Pass@$K$ score significantly surpasses both the base model and Reason-RFT. This clearly highlights Vision-EKIPL’s ability to push the reasoning frontier forward. Its core advantage lies in the integration of high-quality actions generated by external expert models during training. These actions serve as distilled supervision targets, enabling the policy model to learn diverse and novel reasoning strategies that would otherwise be inaccessible through self-sampling.
@@ 5 Conclusion
Appendix A Importance Sampling Correction and Its Properties
This section presents a rigorous derivation and properties of importance sampling when samples are drawn from a general proposal distribution $q(o\mid s)$, rather than the reference policy $\pi_{\theta_{\mathrm{old}}}$. We introduce the unbiased IS transformation, the self-normalized IS (SNIS) estimator, bias bounds, the effective sample size (ESS), and explicit formulas for mixture proposals.
A.1 Basic Identity and Correction
Consider the expectation under the reference policy:
$J\;=\;\mathbb{E}_{o\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid s)}\big[g(o)\big],$
where $g(o)$ denotes any integrable function (in our case, often the product of a policy ratio and an advantage term, e.g., $g(o)=\tfrac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)$). When samples are drawn from a proposal $q(o\mid s)$, the unbiased IS transformation is obtained via the importance weight
$w(o)\;=\;\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)},$
yielding
$J$ $=\mathbb{E}_{o\sim\pi_{\theta_{\mathrm{old}}}}\big[g(o)\big]=\mathbb{E}_{o\sim q}\!\left[\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)}\,g(o)\right]=\mathbb{E}_{o\sim q}\big[w(o)\,g(o)\big].$ (8)
For $g(o)=\dfrac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)$, Eq. (8) reduces to
$J$ $=\mathbb{E}_{o\sim q}\left[\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{q(o\mid s)}\cdot\frac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{\pi_{\theta_{\mathrm{old}}}(o\mid s)}A(o)\right]=\mathbb{E}_{o\sim q}\left[\frac{\pi_{\theta_{\mathrm{new}}}(o\mid s)}{q(o\mid s)}\,A(o)\right].$ (9)
Hence, when using samples from a proposal $q$, the correct unbiased estimator requires the ratio $\tfrac{\pi_{\theta_{\mathrm{new}}}(o)}{q(o)}$, rather than $\tfrac{\pi_{\theta_{\mathrm{new}}}(o)}{\pi_{\theta_{\mathrm{old}}}(o)}$ (unless $q=\pi_{\theta_{\mathrm{old}}}$).
A.2 Sample-Based Estimators
Suppose we draw $o_{1},\dots,o_{n}\sim q$, with weights $w_{i}=w(o_{i})$ and function values $g_{i}=g(o_{i})$.
Unnormalized IS:
$\widehat{J}_{\mathrm{IS}}\;=\;\frac{1}{n}\sum_{i=1}^{n}w_{i}\,g_{i},$
which is unbiased under mild regularity conditions.
Self-Normalized IS (commonly used in practice):
$\widehat{J}_{\mathrm{SNIS}}\;=\;\sum_{i=1}^{n}\tilde{w}_{i}\,g_{i},\qquad\tilde{w}_{i}\;=\;\frac{w_{i}}{\sum_{j=1}^{n}w_{j}}.$
This reduces variance but introduces a small bias that vanishes as $n\to\infty$.
A.3 Effective Sample Size (ESS)
The dispersion of normalized weights can be measured by the effective sample size:
$\mathrm{ESS}\approx\frac{\big(\sum_{i=1}^{n}w_{i}\big)^{2}}{\sum_{i=1}^{n}w_{i}^{2}}.$
When $q$ is close to $\pi_{\theta_{\mathrm{old}}}$, the weights are nearly constant and ESS $\approx n$; if weights are highly imbalanced, ESS becomes much smaller, leading to unstable estimates. In practice, monitoring ESS, weight variance, or high quantiles (e.g., 95% percentile of weights) provides insight into IS quality.
A.4 Bias Bound via Total Variation
If one ignores the correction and directly uses $\mathbb{E}_{q}[g]$ as an approximation to $J$, the bias can be bounded as
$\left|\mathbb{E}_{\pi_{\theta_{\mathrm{old}}}}[g]-\mathbb{E}_{q}[g]\right|$ $=\left|\int g(o)\big(\pi_{\theta_{\mathrm{old}}}(o)-q(o)\big)\,\mathrm{d}o\right|\leq\|g\|_{\infty}\int\big|\pi_{\theta_{\mathrm{old}}}(o)-q(o)\big|\,\mathrm{d}o$
$=2\|g\|_{\infty}\cdot\mathrm{TV}\!\big(\pi_{\theta_{\mathrm{old}}},q\big),$ (10)
where $\mathrm{TV}(p,q)=\tfrac{1}{2}\int|p-q|$ is the total variation distance. Thus, the approximation bias is jointly controlled by the function magnitude and the distributional divergence.
A.5 Explicit Weights for Mixture Proposals
In many settings, the proposal $q$ is a mixture of multiple policies. For instance, if samples are pooled from the reference policy $\pi_{\theta_{\mathrm{old}}}$ and $M$ auxiliary policies $\{\pi_{\phi_{j}}\}_{j=1}^{M}$, the unfiltered proposal can be modeled as
$q(o\mid s)\;=\;\frac{1}{M+1}\Big(\pi_{\theta_{\mathrm{old}}}(o\mid s)+\sum_{j=1}^{M}\pi_{\phi_{j}}(o\mid s)\Big),$
with the importance weight
$w(o)\;=\;\frac{\pi_{\theta_{\mathrm{old}}}(o\mid s)}{\tfrac{1}{M+1}\big(\pi_{\theta_{\mathrm{old}}}(o\mid s)+\sum_{j=1}^{M}\pi_{\phi_{j}}(o\mid s)\big)}.$
If additional selection is performed (e.g., top-$G$ filtering), the actual proposal becomes the conditional distribution
$q_{\text{sel}}(o\mid s)\;=\;\frac{q(o\mid s)\,\mathbf{1}\{o\in S\}}{\int q(o\mid s)\,\mathbf{1}\{o\in S\}\,\mathrm{d}o},$
where $S$ denotes the selection set. Corresponding weights are then computed as $\pi_{\theta_{\mathrm{old}}}(o\mid s)/q_{\text{sel}}(o\mid s)$.
A.6 Algorithmic Pseudocode
Algorithm 1 Importance-Weighted Estimation with Proposal $q$
1:Samples $\{o_{i}\}_{i=1}^{n}\sim q(o\mid s)$; function values $g_{i}=g(o_{i})$; reference and candidate policy densities
2:IS estimate $\widehat{J}_{\mathrm{IS}}$, SNIS estimate $\widehat{J}_{\mathrm{SNIS}}$, ESS
3:Compute weights $w_{i}\leftarrow\dfrac{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid s)}{q(o_{i}\mid s)}$
[…]
5:Normalize: $\tilde{w}_{i}\leftarrow\dfrac{w_{i}}{\sum_{j=1}^{n}w_{j}}$
[…]
8:return $\widehat{J}_{\mathrm{IS}},\ \widehat{J}_{\mathrm{SNIS}},\ \mathrm{ESS}$
A.7 Discussion and Practical Remarks
  • If $q=\pi_{\theta_{\mathrm{old}}}$, then $w\equiv 1$, and Eq. (9) reduces to the standard policy ratio form $\tfrac{\pi_{\theta_{\mathrm{new}}}}{\pi_{\theta_{\mathrm{old}}}}A(o)$.
  • When $q$ deviates substantially from $\pi_{\theta_{\mathrm{old}}}$, IS or SNIS correction is essential. Monitoring ESS and weight statistics provides diagnostics for stability.
  • If $q$ is analytically intractable (e.g., after complex filtering), approximate calculations of $q(o)$ and corresponding bias analysis should be reported. Ablation comparing (A) uncorrected ratios, (B) exact IS, and (C) SNIS can empirically validate the approximation.
A.8 Summary
For samples drawn from a general proposal $q$, unbiased estimation of expectations under $\pi_{\theta_{\mathrm{old}}}$ requires incorporating importance weights $w=\tfrac{\pi_{\theta_{\mathrm{old}}}}{q}$. In the common case $g(o)=\tfrac{\pi_{\theta_{\mathrm{new}}}}{\pi_{\theta_{\mathrm{old}}}}A(o)$, this is equivalent to using the corrected ratio $\tfrac{\pi_{\theta_{\mathrm{new}}}}{q}$. Approximate forms that drop $w$ may still be employed in practice but should be accompanied by theoretical justification (e.g., bias bound in Eq. (10)) and empirical validation via ESS and ablation studies.